# Strong anharmonicity dictates ultralow thermal conductivities of type-I clathrates

### v1

@@ IV Conclusions
In conclusion, we have developed a framework based on VDMFT for the efficient and accurate calculation of thermal transport properties in real materials, including nonperturbative anharmonic effects and both intraband and interband transport mechanisms. We have applied this methodology to understand the temperature-dependent thermal conductivities of type-I clathrate solids. Although we have used a coarse-grained representation of the atomic structure of these type-I clathrates, our results for the thermal conductivity are qualitatively and quantitatively consistent with ab initio calculations. The authors of Ref. 5 calculate a value of 0.97 W/mK for BaGG at 300 K using the BTE-SMA approach with lifetimes extracted from SCP+PT calculations and ignoring interband transport, and the authors of Ref. 35 calculate values of 1.26 W/mK and 0.77 W/mK for BaGG and SrGG, respectively, using a similar approach while also including interband transport. Those works successfully capture the reduction in thermal conductivities of the filled clathrates, even while using conventional PT, because of the atomistic detail in their clathrate structures. In the atomistic configuration (Fig. 1a), the spherical asymmetry of the cage structures enable three-phonon processes that are strictly zero in our spherically symmetric coarse-grained structure (Fig. 1b). However, results with our coarse-grained model do confirm that the mechanism for thermal conductivity reduction in the filled clathrates is due to nonresonant scattering that decreases lifetimes of cage acoustic modes [4, 5, 35], even at frequencies away from those of the guest rattling modes, and that changes in acoustic mode velocity play a relatively minor role.
@@
Finally, as shown in Fig. 4, our calculations are able to capture the experimental high-temperature behavior of BaGG, while the results of Refs. 4, 5 show a steeper scaling with increasing temperature. This discrepancy may be due to role of nonperturbative effects discussed above.
The role of nonperturbative effects has also been suggested by Ikeda et al.[6] as being important for capturing the experimental high-temperature dependence of thermal conductivity. However, Ikeda et al. also discuss that the thermal conductivity peaks at 10 K, a temperature lower than that expected by the traditional Debye model. We calculate thermal conductivities that peak between 25-50 K, but we neglect nuclear quantum effects, which may be important at these low temperatures.
@@
Future work includes applying our VDMFT framework to a fully atomistic description of the type-I clathrate materials to evaluate the relative importance of three-phonon versus multi-phonon scattering processes as well as including quantum nuclear dynamics to understand their effect on thermal conductivities, especially at very low temperatures.
@@
Acknowledgements. This work was supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, under Award No. DESC0023002. We acknowledge computing resources from Columbia University’s Shared Research Computing Facility project, which is supported by NIH Research Facility Improvement Grant No. 1G20RR030893-01, and associated funds from the New York State Empire State Development, Division of Science Technology and Innovation (NYSTAR), Contract No. C090171.

### v2

@@ IV Conclusions
In conclusion, we have developed a framework based on VDMFT for the efficient and accurate calculation of thermal transport properties in real materials, including nonperturbative anharmonic effects and both intraband and interband transport mechanisms. We have applied this methodology to understand the temperature-dependent thermal conductivities of type-I clathrate solids. Although we have used a coarse-grained representation of the atomic structure of these type-I clathrates, our results for the thermal conductivity are qualitatively and quantitatively consistent with ab initio calculations. The authors of Ref. 8 calculate a value of 0.97 W/mK for BaGG at 300 K using the BTE-SMA approach with lifetimes extracted from SCP+PT calculations and ignoring interband transport, and the authors of Ref. 55 calculate values of 1.26 W/mK and 0.77 W/mK for BaGG and SrGG, respectively, using a similar approach while also including interband transport. Those works successfully capture the reduction in thermal conductivities of the filled clathrates, even while using conventional PT, because of the atomistic detail in their clathrate structures (Fig. 1a), which enables three-phonon processes that are strictly zero in our coarse-grained structure (Fig. 1b). Resolving this discrepancy, wherein quite different approaches predict similar conductivities, is an important goal of our future research, which will be enabled by our fully atomistic, first-principles implementation of VDMFT. However, results with our coarse-grained model do confirm that the mechanism for thermal conductivity reduction in the filled clathrates is due to nonresonant scattering that decreases lifetimes of cage acoustic modes [7, 8, 55], even at frequencies away from those of the guest rattling modes, and that changes in acoustic mode velocity play a relatively minor role.
@@
Finally, as shown in Fig. 4, our calculations are able to capture the experimental high-temperature behavior of BaGG, while the results of Refs. 7, 8 show a steeper scaling with increasing temperature. This discrepancy may be due to the role of nonperturbative effects discussed above.
Nonperturbative effects, in the form of a phonon Kondo effect, have also been suggested as being responsible for some unconventional behaviors of the thermal conductivity of type-I clathrates [9], including a narrow maximum at temperatures lower than the Debye temperature and a larger high-temperature conductivity than predicted by perturbation theory. Given the success of DMFT and its associated impurity problem in explaining Kondo physics in electronic systems, our method is uniquely suited to capture any such physics. The hybridization of a dispersionless and dispersive band is indeed suggestive, although the exact expectations of a phonon Kondo effect are unclear, given that phonons lack a Fermi energy and exhibit a trivial maximum in their conductivity (see SM [33] for additional details). Although the Debye temperature of our studied clathrates is 60–100 K, our calculated conductivities exhibit a maximum at lower temperatures of 25–50 K, which are comparable to the Einstein temperature of the rattling modes. Moreover, at high temperature we see a weaker dependence of the conductivity than predicted by perturbation theory. Both of these results are qualitatively in line with experimental observations.
The results of this work can be used to extract design principles for the development of real clathrate materials. The introduction of guest atoms inside clathrate cages increases anharmonicity of the lattice dynamics and lowers the thermal conductivity significantly. Smaller guest atoms, such as Sr compared to Ba, have more free space to move within the cages, resulting in more anharmonic cage-guest interactions as well as lower-frequency X(2) rattling modes with shorter lifetimes due to multiphonon scattering. These factors generally reduce the thermal conductivity, as shown in Fig. 4a. However, these factors can also enable interband transport, especially at higher temperatures, which can elevate the thermal conductivity (Fig. 3a). This competition indicates that a careful trade-off must be considered when designing thermoelectric materials with low thermal conductivities.
@@
In future work, we will apply our VDMFT framework to a fully atomistic description of the type-I clathrate materials to evaluate the relative importance of three-phonon versus multi-phonon scattering processes in real materials with lower symmetry than our coarse-grained model. Moreover, although our calculated conductivity partially captures nuclear quantum effects through its use of Bose-Einstein statistics, future work will study the role of nuclear quantum dynamics, which may be important at very low temperatures.
@@
Acknowledgements. This work was supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, under Award No. DESC0023002. We acknowledge computing resources from Columbia University’s Shared Research Computing Facility project, which is supported by NIH Research Facility Improvement Grant No. 1G20RR030893-01, and associated funds from the New York State Empire State Development, Division of Science Technology and Innovation (NYSTAR), Contract No. C090171 The Flatiron Institute is a division of the Simons Foundation.